please help me solve this:
An isosceles triangle has a perimeter of 82 cm. The lengths of the legs of the triangle are represented by (3x+2) and (5x-14). Find the length of the base of the triangle.

Answer by drk(1908)

In an isosceles triangle, two legs are equal. In this case, we have
(3x+2) = (5x-14). Solve for x.
16 = 2x
8 = x.
Substitute back into the equation to get the legs = 26.
The perimeter = 84 = a + b + c.
84 = a + 26 + 26
84 = a + 52
32 = a.
The base is 32 cm